\begin{thebibliography}{99}

\bibitem{osama}	Osama H. H. Edely, S. A. Mohiuddine, Abdullah K. Noman, Korovkin type approximation theorems obtained through %generalized statistical convergence, \emph{Appl. Math. Lett.} 23(11), 1382-1387 (2010).

\bibitem{LS1} A. D. Gadjiev, C. Orhan,Some approximation theorems by statistical convergence, Rocky Mountain J.math. $32(2002)129-138.$

\bibitem{lp} A. Lupa\c{s}, A $q$-analogue of the Bernstein operator, Seminar on Numerical and Statistical Calculus, \emph{University of Cluj-Napoca}, 9(1987)85-92.

\bibitem{bm} C. Belen, S. A. Mohiuddine Generalized weighted statistical convergence and application. \emph{Applied Mathematics and Computation,} 219(18): 9821-9826 (2013).

\bibitem{inter} George M. Phillips,Interpolation and Approximation by Polynomials, (2003) \emph{Springer}

	\bibitem{pl} George M. Phillips, Bernstein polynomials based on the $q$-integers, The heritage of P.L.Chebyshev, \emph{Ann. Numer. Math.}, 4 (1997) 511-518.

	%
\bibitem{kadak2} Ugur Kadak, On weighted statistical convergence based on $(p,q)$-integers and related approximation theorems for functions of two variables, \emph{Journal of Mathematical Analysis and Applications,} May 2016, DOI: 10.1016/j.jmaa.2016.05.062.\\
			
			
			
\bibitem{kadak5} Ugur Kadak, Weighted statistical convergence based on generalized difference operator involving $(p,q)$-Gamma function and its applications to approximation theorems, Volume 448, Issue 2, (2017), Pages 1633-1650, \emph{Journal of Mathematical Analysis and Applications.}\\
			
\bibitem{kadak6} Ugur Kadak, Vishnu Narayan Mishra and Shikha Pandey,	Chlodowsky type generalization of $(p, q)$-Szasz operators involving Brenke type polynomials, Revista de la Real Academia de Ciencias Exactas, Físicas y Naturales. Serie A, Matemáticas, September 2017, DOI: 10.1007/s13398-017-0439-y.
			
\bibitem{kanat} K. Kanat,  M. Sofyalıoglu,   Some approximation results for Stancu type Lupaş–Schurer operators based on (p, q)-integers, Applied Mathematics and Computation Volume 317, 15 January 2018, Pages 129–142.


\bibitem{king} J. P. King, Positive linear operators which preserve $x^2,$ Acta Math.Hunger. 99(2003)203-208

\bibitem{ogun} Ogun Dogru, Kadir Kanat, On statistical approximation properties of the Kantorovich type Lupa\c{s} operators, Mathematical and Computer Modelling 55 (2012) 1610-1620.
			


\bibitem{khalidjcam} Khalid Khan, D.K. Lobiyal, B$\acute{e}$zier curves based on Lupa\c{s} $(p,q)$-analogue of Bernstein functions in CAGD,  \emph{Journal of Computational and Applied Mathematics} Volume 317, June 2017, Pages 458-477.

\bibitem{mishra2} V.N. Mishra, S. Pandey, On $(p,q)$ Baskakov-Durrmeyer-Stancu Operators, Advances in Applied Clifford Algebras, (2016), DOI: 10.1007/s00006-016-0738-y.
			
\bibitem{mishra3}  V.N. Mishra, S. Pandey; On Chlodowsky variant of $(p, q)$ Kantorovich-Stancu-Schurer operators, International Journal of Analysis and Applications, Vol. 11, No. 1, pp. 28-39 (2016).



\bibitem{mka1}  M. Mursaleen, K. J. Ansari, Asif Khan, On $(p,q)$-analogue of Bernstein Operators, \emph{Applied Mathematics and Computation,} 266 (2015) 874-882, (Erratum:  Appl. Math. Comput. 266 (2015) 874-882.)
			

		
\bibitem{mab} M. Mursaleen and Asif Khan, Statistical Approximation Properties of Modified $q$-Stancu-Beta Operators, Bull. Malays. Math. Sci. Soc. (2) 36(3) (2013), 683–690.
	
\bibitem{mfa}  M. Mursaleen, Faisal Khan and Asif Khan, Statistical approximation for new positive linear operators of Lagrange type,  \emph{Applied Mathematics and Computation,} 232 (2014) 548–558.

\bibitem{bsm}	Naim L. Braha, H. M. Srivastava, S. A. Mohiuddine,  A Korovkin's type approximation theorem for periodic functions via the statistical summability of the generalized de la Vallée Poussin mean.\emph{Applied Mathematics and Computation,} 228, 162-169 (2014).

\bibitem{agritini} O. Agratini, O. Dogru,Weighted statistical approximation by q-szasz type operators that preserve some test functions,Taiwanese j.Math. 14(4)(2010)
	
\bibitem{sn} S.N. Bernstein,Demonstration du theoreme de weierstrass fondee sur la calcul des probabilities, comm.soc.Math.Kharkov 13(1912) 1-2
\bibitem{fast} H. Fast, Sur la convergence statistique, Colloq. Math. 2(1951) 241-244.

\bibitem{wn1} A. Wafi, N. Rao,  Bivariate-Schurer-Stancu operators based on $(p,q)$-integers, general mathematics, 2016 (accepted).
			
			
\bibitem{wn2} A. Wafi, N. Rao, $(p,q)$-Bivariate-Bernstein-Chlowdosky Operators, Filomat, 2016,
(accepted)http://journal.pmf.ni.ac.rs/filomat/index.php/filomat/author/index/completed.

\bibitem{iran} Asif Khan, Vinita Sharma,  Approximation by  $(p,q)$-Lupa\c{s} Stancu Operators Iranian Journal of Mathematical Sciences and Informatics \\
Vol. x, No. x (201x), pp xx-xx

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsfonts}
\usepackage{latexsym,amsmath,amssymb}
\usepackage{graphics,epstopdf}
\usepackage[pdftex]{graphicx}
\usepackage{subcaption}

\bibliography{bibfile}
\newtheorem{definition}{Definition}[section]
\newtheorem{theorem}[definition]{Theorem}

\newtheorem{lemma}[definition]{Lemma}
\numberwithin{equation}{section}

\newtheorem{thm}{Theorem}[section]

\newtheorem{rem}[thm]{Remark}
\numberwithin{equation}{section}

\begin{document}

	\bigskip
	
	\bigskip
	
	\begin{center}
		{\Large \textbf{ On King type modification of $(p,q)$-Lupa\c{s} Bernstein operators.
			}}
			
	\bigskip
			
	\textbf{ Asif Khan, Vinita Sharma }
	and \textbf{Faisal Khan}
			
	Department of\ Mathematics, Aligarh Muslim University,
	Aligarh 202002, India%
	\\[0pt]
			
	akhan.mm@amu.ac.in; vinita.sha23@gmail.com; faisalamu2011@gmail.com
			

	\bigskip
			
	\bigskip

	\textbf{Abstract}
	\end{center}
	\parindent=8mm {\footnotesize {In this paper, a King-type modification of $(p,q)$-Lupa\c{s} Bernstein operators are introduced. The rate of convergence of these operators are studied by  means of modulus of continuity and Lipschitz class functional.\\
	Further, it has been shown that the error estimation of these operators on some subintervals of $[0,1]$ are better than the $(p,q)$-Lupa\c{s} Bernstein operators.\\ 
	\bigskip
			
	{\footnotesize \emph{Keywords and phrases}: $(p,q)$-integers; $(p,q)$-Bernstein operators,   $(p,q)$-Lupa\c{s} Bernstein operators, King type approximation, modulus of continuity.}\\

	{\footnotesize \emph{MSC: primary 65D17; secondary 41A10, 41A25, 41A36.}: \newline

 \section{Introduction}		
First, 	Let us recall certain notations of $(p,q)$-calculus.\\

For any $n\in \mathbb N$, the $(p,q)$-integers are defined as follows:
 	\begin{equation*}
 	\lbrack n]_{p,q}=p^{n-1}+p^{n-2}q+p^{n-3}q^2+...+pq^{n-2}+q^{n-1}\\
 	=\left\{
 	\begin{array}{lll}
 	\frac{p^{n}-q^{n}}{p-q},~~~~~~~~~~~~~~~~\mbox{when $~~p\neq q \neq 1$  } & \\
 	&  \\
 	n~p^{n-1},~~~~~~~~~~~~~~\mbox{ when $p=q\neq1$  } & \\
 	&  \\
 	
 	[n]_q ,~~~~~~~~~~~~~~~~~~~\mbox{when $p=1$  }& \\

 	n ,~~~~~~~~~~~~~~~~~~~~~\mbox{ when $p=q=1.$  }
 	\end{array}%
 	\right.
 	\end{equation*}
 Also the $(p,q)$-binomial coefficient is defined by\\

 $${n \brack k}_{p,q}=\frac{[n]_{p,q}!}{[k]_{p,q}!~[n-k]_{p,q}!}~~\text{for all}~~n, k\in \mathbb N~~\text{with}~~n\ge k.$$
		
When $ p=1$ and $q=1,$ it reduces to the ordinary integers and binomial cofficient respectively.\\

Recently, the applications of $(p,q)$-calculus emerged as a new area in the field of approximation theory. The $(p,q)$-calculus development has led to the discovery of various
			generalizations of Bernstein polynomials based on $(p,q)$-integers. The purpose of
			these generalizations is to provide appropriate and powerful tools to
			application areas such as computer-aided geometric
			design, numerical analysis,  and solutions of differential equations.\\
			
			Mursaleen \textit{et al}  \cite{mka1} first introduced $(p,q)$-calculus in approximation theory and constructed the {\ $(p,q)$-analogue of Bernstein operators defined as follows  for $0<q<p\leq 1$:\\

\begin{equation}\label{ee1}
B_{n,p,q}(f;x)=\frac1{p^{\frac{n(n-1)}2}}\sum\limits_{k=0}^{n}\left[
\begin{array}{c}
n \\
k%
\end{array}%
\right]_{p,q}p^{\frac{k(k-1)}{2}}x^{k}\prod \limits_{s=0}^{n-k-1}(p^{s}-q^{s}x)~~f\left( \frac{
[k]_{p,q}}{p^{k-n}[n]_{p,q}}\right) ,~~x\in \lbrack [0,1].
\end{equation}

			where

				\begin{align*}
				(1-x)^{n}_{p,q}&=\prod\limits_{s=0}^{n-1}(p^s-q^{s}x) =(1-x)(p-qx)(p^{2}-q^{2}x)...(p^{n-1}-q^{n-1}x)\\
				&=\sum\limits_{k=0}^{n} {(-1)}^{k}p^{\frac{(n-k)(n-k-1)}{2}} q^{\frac{k(k-1)}{2}}\left[
				\begin{array}{c}
				n \\
				k
				\end{array}%
				\right] _{p,q}x^{k}.
				\end{align*}

					Note when $p=1,$ $(p,q)$-Bernstein Operators given by $(\ref{ee1})$ turns out to be Phillips $q$-Bernstein Operators \cite{pl}.\\

The $q$-analogue of Bernstein operators \cite{sn} introduced by Lupa\c{s} \cite{lp} are as follows:\\

	\begin{equation}\label{e1.1}
	L_{n}(f;p;q;x)=\sum\limits_{k=0}^{n}f\biggl{(}\frac{k}{n}\biggl{)}b_{nk}(q;x),~~f\in C[0,1]~~ and~~  x\in [0,1].
	\end{equation}
	where\\
	
\begin{equation}\label{15e}
b_{nk}(q;x)=\frac{\left[\begin{array}{c}
	n \\
	k%
	\end{array}
	\right] _{q} q^{\frac{k(k-1)}{2}}~x^{k}~(1-x)^{n-k}}{\prod\limits_{j=0}^{n-1}\{(1-x)+q^{j} x\}}.
\end{equation}

Recently, Khalid et al. defined  $(p,q)$-analogue of Lupa\c{s} Bernstein operators \cite{khalidjcam}  as follows :\\

For any $p>0$ and $q>0,$ the linear operators $L^{n}_{p,q}:$ $C[0,1] \rightarrow C[0,1]$
\begin{equation}\label{e9a}
L^{n}_{p,q}(f;x)=L_{n}(f;p;q;x)= \sum\limits_{k=0}^{n}~~ \frac{f \bigg(\frac{p^{n-k}~[k]_{p,q}}{[n]_{p,q}}\bigg) ~\left[\begin{array}{c}
	n \\
	k%
	\end{array}%
	\right] _{p,q} p^{\frac{(n-k)(n-k-1)}{2}} q^{\frac{k(k-1)}{2}}~x^{k}~(1-x)^{n-k}}{\prod\limits_{j=1}^{n}\{p^{j-1}(1-x)+q^{j-1} x\}}.
\end{equation}

We recall the following lemma for the above operators.\\

\begin{lemma}\label{b1}\cite{khalidjcam}
	The following equalities are true\\
	
(1.)	$L_{n}(1;p;q;x)= 1$\\

(2.)	$L_{n}(t;p;q;x)= x$\\
	
(3.)	$L_{n}(t^{2};p;q;x)= x^2 +\frac{x(1-x)p^{n-1}}{[n]_{p,q}}~-~\frac{x^2(p-q)(1-x)}{p(1-x+qx}\big(1 - \frac{p^{n-1}}{[n]_{p,q}})$\\
\end{lemma}

Here, we know that the opeator $L_{n}(t^{2};p;q;x)$ do not preserve the test function $e_{2}$.\\

 In 2003, King \cite{king} introduced a non- trivial sequence of opeators preserving the functions $e_0$ and $e_2$ where $(e_i~=~x^i,i~=~0,1,2).$

He also proved that these operators have a better rate of convergence than the classical Bernstein polynomials whenever $0 \leq x \leq \frac{1}{3}$. In \cite{agritini}, Agratini and dogru introduced a King type modification of $q$-Szasz-Mirakjan type operators and they proved that their operators have a better rate of convergence than the classical
-Szasz-Mirakjan operators.\\

One can refer \cite{ kadak2,kadak5,kadak6,kanat,ogun,mishra2,mishra3,wn2} for similar recent works  based on  $(p,q)$-integers in the field of approximation theory. \\

In this paper, we consider  $0<q<p \leq 1$ and a King type modification of Lupa\c{s} Bernstein operators defined in \cite{lp} and investigate the statistical approximation properties of these operators. At last we show that this type of modification gives us better error estimation on some subintervals of $[0,1]$ than the classical  $(p,q)$-Lupa\c{s} Bernsrein operators. In case $p=1,$ it reduces to King type modification of $q$-Lupa\c{s} Bernstein operators. \\

\section{Construction of Operators}
 Now, we construct the King type modification of $(p,q)$-Lupa\c{s} Bernsrein operators (\ref{e9a}) which preserve monomials $e_i(x)= x^{i}$ for $(i= 0,2)$. For this study, we consider $0<q<p\leq 1$  satisfying the following condition\\
 \begin{equation}\label{bb}
 pq([n]-1) >p^{n}(p-q)
 \end{equation}
  for $n\geq 2.$\\

Let $r_{n}(x)$ be a sequence of real valued continuous functions defined on $[0,1]$ with $0\leq r_{n}(x)\leq1$. Let us consider the following operators:\\
		
\begin{equation}\label{18e}
L^*_{n}(f;p;q;x)= \sum\limits_{k=0}^{n}~~ \frac{f \bigg(\frac{p^{n-k}~[k]_{p,q}}{[n]_{p,q}}\bigg) ~\left[\begin{array}{c}
	n \\
	k%
	\end{array}%
	\right] _{p,q} p^{\frac{(n-k)(n-k-1)}{2}} q^{\frac{k(k-1)}{2}}~r_{n}(x)^{k}~(1-r_{n}(x))^{n-k}}{\prod\limits_{j=0}^{n-1}\{p^{j}(1-r_{n}(x))+q^{j} r_{n}(x)},
\end{equation}		
		
Where $f \in C[0,1] $, $ x \in [0,1]$ and $n\in N \backslash {0,1}$. it is clear that the operator $L^{*}_{n}(f,p,q;x)$ are positive and linear. Observe that, if we choose $r_{n}(x)=x$ then it turn out to be $(p,q)$-Lupa\c{s} Bernstein operators.

\begin{lemma}\label{b2}
	$L^*_{n}(f;p;q;x)$ satisfy the following properties.\\
	(1.)$L^*_{n}(e_0;p;q;x) = 1$\\
	(2.)$L^*_{n}(e_1;p;q;x) = r_{n}(x)$\\
	(3.)$L^*_{n}(f e_{2};p;q;x) = r_{n}^2(x) +\frac{x(1-x)p^{n-1}}{[n]_{p,q}}~-~\frac{r_{n}(x)^2(p-q)(1-r_{n}(x)}{p(1-r_{n}(x)+qr_{n}(x)}\big(1~-~\frac{p^{n-1}}{[n]_{p,q}}).$\\
\end{lemma}
\textbf{Note:} For our convenience, we denote $[n]_{p,q}~=~[n].$\\
		
Under the condition ( \ref{bb}), if we take\\

\begin{equation}\label{b3}
r_{n}(x)~=~-\frac{p^n+x^2[n](p-q)}{2(p^{n-1}q-p^{n}+q^2[n-1])}~+~\frac{\sqrt{{p^{2n}+x^{4}[n]^{2}(p-q)^{2}+2x^{2}[n](2pq([n]-1)-p^n(p-q))}}}{2(p^{n-1}q-p^{n}+q^2[n-1])},
\end{equation}		

then $L^*_{n}(f;p;q;x)$	preserve monomials, $L^*_{n}(e_0;p;q;x) = e_0(x) = 1$ and $L^*_{n}(e_2;p;q;x) = e_2 = x^2,$ for $n \geq 2$ .\\
Also $0 \leq r_{n}(x)\leq 1$ for $r_{n}(x)$ defined in (\ref{b3}).\\

From (\ref{bb}), we have\\
 \begin{equation}\label{cc}
2pq([n]-1)~-~p^{n}(p-q) > p^{n}(p-q)
 \end{equation}

 Using the inequality (\ref{cc}) we get\\

  \begin{equation}\label{c1}
p^{2n}~+~2x^{2}[n]p^{n}(p-q)~+~x^{4}[n]^{2}(p-q)^{2} = \big(p^{n}~+~x^{2}[n](p-q)\big)^{2}.
  \end{equation}

From above equality, we get $r_{n}(x)\geq 0$
 under the condition (\ref{bb}), since $(1-x)^2 \geq 0$ for $0 \leq x \leq 1,$ we have\\

   \begin{equation}\label{g1}
   p^{2n}~+~2x^{2}[n](2pq([n]-1)-p^{n}(p-q))~+~x^{4}[n]^{2}(p-q)^{2} \leq  \big(2(p^{n-1}q-p^{n}+q^2[n-1])+p^{n}~+~x^{2}[n](p-q)\big)^{2}.
   \end{equation}
    If we use (\ref{g1}) in (\ref{b3}) then we get $r_n(x)\leq 1$.\\
 		
\begin{rem}\label{r5.1}
	For $q\in(0,1)$ and $p\in(q,1]$, it is obvious that
	$\lim\limits_{n\to\infty}[n]_{p,q}=0 $ or $\frac1{p-q}$. In order to reach to convergence
	results of the operator $L^*_{n}(f;p;q;x),$ we take a sequence $q_n\in(0,1)$ and $p_n\in(q_n,1]$
	such that $\lim\limits_{n\to\infty}p_n=a,$ $\lim\limits_{n\to\infty}q_n=1$ and $\lim\limits_{n\to\infty}p_n^n=1,$ $\lim\limits_{n\to\infty}q_n^n=1$. So we get
	$\lim\limits_{n\to\infty}[n]_{p_n,q_n}=\infty$.
\end{rem}

\begin{thm}\label{1h1}
	Let $L^*_{n}(f;p;q;x)$ be the sequence of operators and the sequence $ p=p_n$ and $q=q_n$ satisfying Remark $\ref{r5.1}$  then for any function $f\in C[0,1]$\\
	$$\lim\limits_{n\to\infty}\lvert L^*_{n}(f;p;q;x_0)-f(x_0)\rvert =0$$\\
	for fixed $x_0\in [0,1]$\\
\end{thm}

\section{The Rates of Convergence}
		
The modulus of continuity for the space of function $ f\in C[0,1]$  is defined by\\

$$ w(f;\delta)=\sup\limits_{x,t\in C[0,1],~~ |t-x|<\delta} |f(t)-f(x)|$$\\

where ${w}(f;\delta)$ satisfies the following conditions:~~for all $f\in C[0,1],$\\

\begin{equation}\label{e118}
\lim\limits_{\delta\rightarrow0}~w (f;\delta) = 0.
\end{equation}
and
\begin{equation}\label{e119}
|f(t)-f(x)|\leq w(f;\delta)\bigg(\frac{|t-x|}{\delta}+ 1\bigg)		
\end{equation}				
		
Recall that, In \cite{iran} we obtained the following rate of convergence for the operators $(\ref{e1.1})$ for every $ f\in C[0,1]$ and $\delta> 0.$\\

\begin{equation}\label{e110}
|L_{n}(f;p;q;x)-f(x)|\leq w(f;\delta)\bigg(\frac{1}{\delta} \sqrt{{\frac{x(1-x)}{[n]}}}+ 1\bigg)		
\end{equation}

Now, we compute the rates of convergence of the operators $L^*_{n}(f;p;q;x)$ given by $(\ref{18e})$ to $f(x)$  by means of the modulus of continuity. and we also show that our error estimation is better than the $(p,q)$-Lupa\c{s} operator given by (\ref{e9a}).\\

\begin{theorem}\label{tt}
	Let $(p_n)$ and $(q_n)$ are the sequences satisfying remark (\ref{r5.1}) for each $n \geq 2$. For fixed $x \in [0,1],$ $f\in C[0,1]$ and $\delta_{n} >0$, we have\\
	
	$$|L^*_{n}(f,p_n,q_n;x)-f(x)|\leq 2w(f;\delta_{n}(x))$$\\
	Where\\
	
\begin{equation}
\delta_{n}(x)~=~  \sqrt {2x^{2} ~+~x\bigg(\frac{p^{n}_n+x^2[n](p_n-q_n)}{2(p^{n-1}_nq_n-p^{n}_n+q^{2}_n[n-1])}~-~\frac{\sqrt{{p^{2n}_n+x^{4}[n]^{2}(p_n-q_n)^{2}+2x^{2}[n](2p_nq_n([n]-1)-p^{n}_n(p_n-q_n))}}}{2(p^{n-1}_nq_n-p^{n}_n+q^{2}_n[n-1])}\bigg)}\tiny\\
\end{equation}
		
\end{theorem}

\begin{theorem}\label{2h2}
For all $f \in Lip_M(\rho)$\\

$\|L_{n}(f,p_n,q_n;x)-f(x)\|_{C[0,1]}~\leq~M\delta^{\rho}_n(x)$\\
		
where\\

$\delta_n(x) = \sqrt{\frac{x(1-x)}{[n]}}$
		
and M is a positive constant.
\end{theorem}

\begin{theorem}\label{3h3}
	For all $f \in Lip_M(\rho)$, under the condition (\ref{bb}) for $p~=~p_n$ and $q~=~q_n$, we have\\
	
$\|L^*_{n}(f,p_n,q_n;x)-f(x)\|_{C[0,1]}~\leq~M\delta^{\rho}_n(x)$\\
		
where\\
		
$\delta_n(x) =\sqrt{ 2x^2~+~x\bigg(\frac{p^n_n+x^2[n](p_n-q_n)}{2(p^{n-1}_nq_n-p^{n}_n+q^2_n[n-1])}~-~\frac{\surd{p^{2n}_n+x^{4}[n]^{2}(p_n-q_n)^{2}+2x^{2}[n](p^n(p_n-q_n))}}{2(p^{n-1}_nq_n-p^{n}_n+q^2_n[n-1])}\bigg)}$
		
and M is a positive constant.
\end{theorem}

\section{The Rates of Statistical Convergence}

	At this point, let us recall the concept of statistical convergence.
	The statistical convergence which was introduced by Fast \cite{LS1} in 1951, is an important research area in approximation theory. In \cite{LS1}, Gadjiev and Orhan used the concept of statistical convergence in approximation theory. They proved a Bohman-Korovkin type theorem for statistical convergence.\\
	Recently, statistical approximation properties of many operators are investigated \cite{bm,mab,mfa,bsm}.\\

	A sequence $x = (x_k)$ is said to be statistically convergent to a number L if for every $\epsilon > 0,$\\
	
	$$\delta \{K\in\textbf{ N}:|x_k -L|\geq \varepsilon\}~=~0,$$\\
	
	where $\delta(K)$ is the natural density of the set $K \subseteq \textbf{ N}.$\\
	
	The density of subset $K\subseteq N$ is defined by\\
	
	$$\delta(K):= \lim\limits_n\frac{1}{n}\{\text{the  number}~~ k \leq n : k\in K\}$$\\
	
	whenever the limit exists.\\

	For instance, $\delta(\textbf{N})~=~1, ~~\delta\{2K:k\in\textbf{ N}\}~=~\frac{1}{2}~~ \text{and}  ~~\delta\{k^2:K\in\textbf{ N}\}~=~0.$ \\
	
	To emphasize the importance of the statistical convergence, we have an example: The sequence\\
	\begin{equation}
	X_k =
	\left\{
	\begin{array}{lll}
	{L_1;~~~~~ if~~k = m^2}, & \\
	&\\
	{L_2;~~~~~~if~~k \neq m^2}. &  \\
	&\\
	\end{array}%
	\right.
	where~~ m\in\textbf{ N}
	\end{equation}
	
	is statistically convergent to $L_2$ but not convergent in ordinary sense when $L_1 \neq L_2.$ We note that any convergent sequence is statiscally convergent but not conversely. \\
	
	Now we consider sequences $ q = q_n$ and $p =p_n$ such that:\\
	
	\begin{equation}\label{1e1}
	st-\lim\limits_nq_n = 1,~~~~~~~st-\lim\limits_np_{n} = 1,~~~~~~~~ st-\lim\limits_nq^{n}_n = 1~~~~~~ \text{and}~~~~~~ st-\lim\limits_np^{n}_{n} = 1.
	\end{equation}
	
\begin{theorem}\label{4h4}
	If the sequences $p = p_n$ and $q= q_n$ satisfies the condition given in (\ref{1e1}), then\\
\begin{equation}\label{22}
|L^{*}_{n}(f,p_n.q_n;x)-f(x)|\leq 2w(f;\surd \delta_{n,x})\\
\end{equation}

for all $f \in C[0,1]$, where\\

$\delta_{n,x}~=~ L^{*}_{n}((t-x),^2p_n.q_n;x)$
\end{theorem}

\begin{theorem}\label{5h5}
If the sequences $p = p_n$ and $q= q_n$ satisfies the condition given in (\ref{1e1}), if $f\in C[0,1]$ then\\
	
$\|L_{n}(f,p_n,q_n;x)-f(x)\|_{C[0,1]}~\leq~2w(f;\delta_n)$\\

where\\
\begin{equation}\label{p6}
\delta_n~=~ \sqrt{\frac{2}{9[n]}}\\
\end{equation}

\end{theorem}

\begin{theorem}\label{6h6}
If the sequences $p = p_n$ and $q= q_n$ satisfies the condition given in (\ref{1e1}), if $f\in Lip_M(\rho)$ then\\

\begin{equation}\label{r2}
|L^{*}_{n}(f,p_n,q_n;x)-f(x)|~\leq~M\delta^{\rho}_n(x)\\
\end{equation}

where\\

$\delta_n(x) = \sqrt{\frac{2}{9[n]}}$\\

\end{theorem}


\end{document}